OpenStudy (anonymous):

he atomic radius of metal X is 1.20 × 102 picometers (pm) and a crystal of metal X has a unit cell that is face-centered cubic. Calculate the density of metal X (atomic weight = 42.3 g/mol)

OpenStudy (aaronq):

start with the density formula: \(\rho=\dfrac{m}{V}\) in FCC there are 4 atoms per unit cell: \(\rho=\dfrac{4*m_{atom}}{V}\) for FCC the edge of the unit cell is \(a=r\sqrt{8}\), so: \(\rho=\dfrac{4*m_{atom}}{(r\sqrt{8})^3}\) last the mass of 1 atom is found by dividing the molar mass by \(N_A\) (avogadros number) \(\rho=\dfrac{4*\dfrac{2.3\;g/mol}{N_A}}{(r\sqrt{8})^3}\) now you just need to convert the radius to units you want before plugging it in

OpenStudy (anonymous):

@Luigi0210 could you help me out here?

OpenStudy (deadpoolroxz):

that haz the answer

OpenStudy (anonymous):

my question has one of the numbers different, the radius is 125pm

OpenStudy (hwyl):

stop cheating @deadpoolroxz

OpenStudy (anonymous):

i get how to do it and i've done the math 4 times now, i cant tell what i'm doing wrong